OpenStudy (hannahc234):

ABC has vertices A(1,5), B(9,7),and C (7.13). M is the midpoint of AB. And N is the midpoint of BC. Which of the following is the length of MN? A. 3 B. 4 C. 5 D. 6

OpenStudy (hannahc234):

@Wholock

OpenStudy (wholock):

Okay so here you will have to find the midpoint of AB and BC, and then use the distance formula to find the length of the segment MN.

OpenStudy (wholock):

For AB, you have (1 + 9) / 2 =? and (5 + 7) / 2 =?

OpenStudy (wholock):

That gives you the midpoint, or point M of the segment.

OpenStudy (wholock):

Now for BC (9 + 7) / 2 =? and (7 + 13) / 2 =? And now you have point N.

OpenStudy (wholock):

Got the points?

OpenStudy (hannahc234):

Yes

OpenStudy (wholock):

You should have M(5, 6) and N(8, 10) Plug these into the distance formula d = sqrt((x2 - x1)^2 + (y2 - y1)^2)

OpenStudy (hannahc234):

I got 5

OpenStudy (wholock):

That's right!

OpenStudy (hannahc234):

Yay! Thank you so much! You're the best.

OpenStudy (wholock):

No problem, anytime!

OpenStudy (hannahc234):

I'll definitely need more help soon thanks so much.

OpenStudy (wholock):

Okay! Just let me know!